OpenStudy (anonymous):

Draw the graphs of 2x-y-1 = 0 and 2x + y = 9; Write down the co-ordinates of the point of intersection of the two lines.

OpenStudy (anonymous):

y=2x-1 y= 9-2x 2x-1=9-2x 4x=10 point of intersection: x=5/2

OpenStudy (anonymous):

2x + y = 9 2x - y = 1 x = 2.5 y = 4 at ( 2.5 , 4 )

OpenStudy (anonymous):

co ordinates is x and y co ordinate